# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

### v2

@@ 4.6.1 Parallel Processing on the GPU
    SRO Calculation: Threads compute the SRO for their assigned atomic configurations using low-precision formats (FP16, Uint16) to reduce data footprint and memory access latency, while maintaining sufficient accuracy.
@@
    Lattice Structure Storage: The lattice structure [41] is stored in shared memory to minimize memory access time during each MC iteration. This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
4.6.2 Iterative Fitness Calculation
Each thread in the block computes the fitness of the perturbed solution during every MC iteration. The fitness, derived from the SRO values calculated in registers, is updated following each perturbation:
$F_{\text{new}}=F_{\text{prev}}+\Delta F_{\text{perturb}},$ (5)
where $\Delta F_{\text{perturb}}$ represents the fitness change from the most recent atomic position perturbation. This value is calculated using the incremental SRO update method.
4.6.3 Fitness Sorting and Selection
After computing the updated fitness in registers, threads within the block collaborate to sort their fitness values using shared memory. The sorting process ensures the rapid identification of the best-performing configuration, which updates the current optimal solution stored in shared memory:
$F_{\text{best}}=\min\left(F_{\text{new}}\right)$ (6)
This guarantees that each block retains the best solution found during the current iteration.
4.6.4 Convergence Check and Termination
The MC search continues iterating until convergence is achieved. Convergence is defined as the point where the fitness of the current best solution shows no significant improvement over a fixed number of steps. Specifically, if $\Delta F_{\text{best}}=0$ for a predefined threshold of iterations, the threads within the block terminate the search. The final result, including the atomic configuration and the associated fitness, is then written back to global GPU memory for subsequent analysis.
This GPU-based approach significantly accelerates the MC search by distributing tasks across multiple threads and blocks, enabling the efficient handling of large-scale systems.
@@ 5 Acknowledgment
This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits,, Peking University. The author acknowledges the financial support provided by the National Natural Science Foundation of China (Grant No. 61874006).
@@ Appendix A Supplementary Data
Table 2: Performance comparison between PyHEA and existing methods under parallel computing conditions. Both PyHEA and SCRAPs utilized 24 CPU cores (Intel Xeon Gold 6132 @ 2.60GHz) for parallel execution, with additional GPU (single NVIDIA 4090) results shown for PyHEA. The table presents computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures of varying sizes (#At) and species numbers (#Sp). All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks (depth 10), and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc). Short-range order was optimized over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
  Type #Spa #Atb ATAT SCRAPs PyHEA PyHEA Speedupc
Parallel Parallel GPU CPU GPU
@@
    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
@@
    ATAT data from Ref. [23]. All times in minutes.

### v3

@@ 4.6.4 Convergence Criteria
    The fitness value $F_{\text{best}}$ remains unchanged for a number of iterations exceeding the search depth threshold defined in subsection 4.2, or
@@
    The maximum number of Monte Carlo steps is reached.
Upon convergence, the final atomic configuration and its associated fitness value are transferred from GPU to CPU memory for subsequent analysis or further refinement cycles. This GPU-optimized implementation achieves significant speedups over CPU-based calculations (see Table 1), enabling the efficient handling of systems containing hundreds of thousands of atoms while maintaining the accuracy of SRO optimization. The approach seamlessly integrates with the broader optimization framework described in section 4, providing a scalable solution for large-scale HEA modeling.
4.7 Molecular Dynamics Simulations
Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40] for two distinct purposes: generating configurations enriched with SRO features and validating the large-scale structures generated by PyHEA.
To generate configurations with target SRO properties, MD simulations were conducted on a 500-atom system under specific temperature and pressure conditions. An NPT ensemble with a timestep of 2 fs was used, and the simulation ran for 200 ps. To accelerate SRO formation, an MC-MD hybrid method was applied, involving random atom swaps every 0.2 ps. The final configuration was extracted for subsequent analysis, ensuring alignment with the desired SRO characteristics.
For validation, MD simulations were carried out on PyHEA-generated configurations containing up to 256,000 atoms. These simulations, also performed in an NPT ensemble with a 2 fs timestep, ran for 10 ps. RDFs were calculated to evaluate the structural reliability of the large-scale configurations, comparing them against known computational and experimental benchmarks.These complementary simulations ensured that SRO structures were accurately generated in small systems and validated in large-scale configurations, providing a robust framework for modeling high-entropy alloys.
Overall, these integrated computational techniques—global and local searches, incremental SRO updates, GPU acceleration, and complementary MD—form a robust toolkit for generating and analyzing large-scale HEA models efficiently and accurately.
@@ 5 Acknowledgment
This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits, Peking University. The author acknowledges the financial support provided by the National Natural Science Foundation of China (Grant No. 61874006).
@@ Appendix A Supplementary Data
Table 2: Performance comparison between PyHEA and existing methods under parallel computing conditions. Both PyHEA and SCRAPs utilized 24 CPU cores (Intel Xeon Gold 6132 @ 2.60GHz) for parallel execution, with additional GPU (single NVIDIA 4090) results shown for PyHEA. The table presents computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures of varying sizes (#At) and species numbers (#Sp). All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks, a search depth of 10, and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc). SRO was optimized over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT-mcsqs in brackets].
  Type #Spa #Atb ATAT- SCRAPs PyHEA PyHEA Speedupc
mcsqs Parallel Parallel GPU CPU GPU
@@
    Speedup ratios: value vs SCRAPs [value vs ATAT-mcsqs in brackets].
@@
    ATAT-mcsqs data from Ref. [23]. All times in minutes.